OpenStudy (anonymous):

the currency in argentina is the peso.the exchange rate is approximately $3=1 peso. at this rate how many pesos would you get if you exchanged $121.10

OpenStudy (anonymous):

does any one have and answer?

OpenStudy (anonymous):

40.36 Pesos I think

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

Lets see $3-1 Peso so $121.10 / 3 = 40.36 I am 80 % sure :D is it multiple choice?

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

3/1 = 121.10/x (cross multiply) (3)(x) = (1)(121.10) 3x = 121.10 x = 121.10/3 = 40.36 I agree with TheMathDude :)

OpenStudy (anonymous):

:D